OpenStudy (anonymous):

Give the 5th and 10th term for the sequence formula below. tn = n + (n+1) + (n+2)

OpenStudy (anonymous):

a= n d=1 t5 = n + (5-1)1 t10 = n + (10 - 1)1

OpenStudy (anonymous):

Just Sub n =10 and n= 5 in the eq